OpenStudy (anonymous):

Find the surface area of a sphere that has a volume of 27 m3

OpenStudy (unklerhaukus):

\[SA_{\text{sphere}}=4\pi r^3\] \[V_\text{sphere}=\tfrac43\pi r^2\]

OpenStudy (unklerhaukus):

look at the second equation first, solve for \(r\) the radius